# FRAP: Faithful and Realistic Text-to-Image Generation with Adaptive Prompt Weighting

### v1

@@ Abstract
Text-to-image (T2I) diffusion models have demonstrated impressive capabilities in generating high-quality images given a text prompt. However, ensuring the prompt-image alignment remains a considerable challenge, i.e., generating images that faithfully align with the prompt’s semantics. Recent works attempt to improve the faithfulness by optimizing the latent code, which potentially could cause the latent code to go out-of-distribution and thus produce unrealistic images. In this paper, we propose FRAP, a simple, yet effective approach based on adaptively adjusting the per-token prompt weights to improve prompt-image alignment and authenticity of the generated images. We design an online algorithm to adaptively update each token’s weight coefficient, which is achieved by minimizing a unified objective function that encourages object presence and the binding of object-modifier pairs. Through extensive evaluations, we show FRAP generates images with significantly higher prompt-image alignment to prompts from complex datasets, while having a lower average latency compared to recent latent code optimization methods, e.g., $4$ seconds faster than D&B on the COCO-Subject dataset. Furthermore, through visual comparisons and evaluation on the CLIP-IQA-Real metric, we show that FRAP not only improves prompt-image alignment but also generates more authentic images with realistic appearances. We also explore combining FRAP with prompt rewriting LLM to recover their degraded prompt-image alignment, where we observe improvements in both prompt-image alignment and image quality.
Introduction
@@
Recent text-to-image (T2I) diffusion models (Rombach et al. 2022; Chang et al. 2023; Kang et al. 2023) have demonstrated impressive capabilities in synthesizing photo-realistic images given a text prompt. Faithfulness and authenticity are among the most important aspects to assess the quality of AI generated content, where faithfulness is evaluated by prompt-image alignment metrics (Radford et al. 2021; Li et al. 2022; Chefer et al. 2023) and authenticity reflects how realistic the image appears and is usually evaluated by image quality assessment metrics (Wang, Chan, and Loy 2023; Wu et al. 2023). While recent T2I diffusion models are capable of generating realistic images with high aesthetic quality, it has been shown (Chefer et al. 2023; Feng et al. 2022; Wang et al. 2022) that the AI generated images may not always faithfully align with the semantics of the given text prompt. One typical issue that compromises the model’s ability to align with the given prompt is catastrophic neglect (Chefer et al. 2023), where one or more objects are missing in the generated image. Another common alignment issue is incorrect modifier binding (Li et al. 2023), where the modifiers, e.g., color, texture, etc., are either bound to the wrong object mentioned in the prompt or ignored. The generation can also result in an unrealistic mix or hybrid of different objects.
@@
To improve prompt-image alignment, Attend & Excite (A&E) (Chefer et al. 2023) introduces the Generative Semantic Nursing (GSN) method, which iteratively adjusts the latent code throughout the reverse generative process (RGP) by optimizing a loss function defined on cross-attention (CA) maps to improve attention to under-attended objects. Similarly, Divide & Bind (D&B) (Li et al. 2023) uses GSN to improve the binding between objects and modifiers, by designing an optimization objective that aligns the attention of objects and their related modifiers. Although GSN can alleviate the semantic alignment issues, yet the iterative updates to latent code could potentially drive the latent code out-of-distribution (OOD), resulting in less realistic images being generated. In addition, several recent works investigate prompt optimization (Hao et al. 2023; Valerio et al. 2023; Mañas et al. 2024), which aims to improve image aesthetics while trying to maintain the prompt-image alignment, by rewriting the user-provided text prompt with a large language model (LLM). Despite improvements on image aesthetic quality, Mañas et al. (2024) show that rewriting prompts primarily for aesthetics enhancement could harm the semantic alignment between the image and prompt. Therefore, it remains a significant challenge to generate images to ensure prompt-image alignment while achieving realistic appearance and high aesthetic quality.
@@
In this paper, we introduce Faithful and Realistic Text-to-Image Generation with Adaptive Prompt Weighting (FRAP), which adaptively adjusts the per-token prompt weighting throughout the RGP of a diffusion model to improve the prompt-image alignment while still generating high-quality images with realistic appearance. Unlike the conventional manual prompt weighting techniques (Face 2023a; Stewart 2023), we design an online optimization algorithm which is performed during inference and updates each token’s weight coefficient in the prompt to adaptively emphasize or de-emphasize certain tokens at different time-steps, by minimizing a unified objective function defined based on the cross-attention maps to strengthen the presence of the prompted object and encourage modifier binding to objects. Since we optimize each prompt token’s weight coefficient, our approach avoids the above-mentioned out-of-distribution issue associated with directly modifying the latent code. Furthermore, we also introduce a novel usage of FRAP by combining it with LLM-based prompt optimizer to achieve outstanding enhancement to both prompt alignment and generation quality.
@@
We extensively evaluate the faithfulness, authenticity and aesthetic quality of FRAP-generated images via prompt-image alignment metrics and a range of image quality assessment metrics and make the following observations:
@@
  • We observe that FRAP achieves significantly higher faithfulness than recent methods on all prompt-image alignment metrics on the Color-Obj-Scene, COCO-Subject, and COCO-Attribute datasets (Li et al. 2023) which contain complex and challenging prompts, while remaining on par with these methods on simple datasets with prompts created from templates. In the meantime, the images generated by FRAP demonstrate higher authenticity with more realistic appearance in terms of the CLIP-IQA-Real measure (Wang, Chan, and Loy 2023).
@@
  • While improving prompt-image alignment and generating more realistic and authentic images, our method requires lower average inference latency (e.g., $4$ seconds faster than D&B on the COCO-Subject dataset) and lower number of UNet calls by not making repeated calls of UNet at each time-step during inference, in contrast to the recent GSN methods (Chefer et al. 2023; Li et al. 2023) which calls the UNet multiple times in a refinement loop at each time-step, especially on datasets with complex prompts that require many repeated calls.
@@
  • We observe that although the LLM prompt rewriting method Promptist (Hao et al. 2023) improves the image aesthetic quality, it could harm the semantic alignment between the image and prompt. Moreover, our adaptive prompt weighting approach can be easily integrated with recent prompt rewrite methods (Hao et al. 2023; Valerio et al. 2023; Mañas et al. 2024). We explore applying FRAP to the rewritten prompts of Promptist to recover their degraded prompt-image alignment, where we observe improvements in both the prompt-image alignment and image quality.
Preliminaries
Stable Diffusion (SD)
@@
We implement our approach based on the open-source SOTA T2I latent diffusion model SD (Rombach et al. 2022). Unlike traditional diffusion models (such as (Sohl-Dickstein et al. 2015; Ho, Jain, and Abbeel 2020; Song, Meng, and Ermon 2020)) that work in the pixel space, SD operates in an autoencoder’s latent space, which significantly reduces the amount of compute resources required. SD is composed of several components, including a Variational Auto-Encoder (VAE) (Kingma and Welling 2013; Rezende, Mohamed, and Wierstra 2014)111 The VAE encoder $\mathcal{E}$ maps an input image $x$ into its spatial latent representation $z=\mathcal{E}(x)$, while the decoder $\mathcal{D}$ attempts to reconstruct the same image from $z$ such that $\mathcal{D}\big{(}\mathcal{E}(x)\big{)}\approx x$. , a UNet (Ronneberger, Fischer, and Brox 2015) that operates in the latent space and serves as the backbone for image generation, and a Text Encoder that is used to obtain embeddings from the text prompt.
UNet.
@@
For training the UNet, the forward diffusion process (Ho, Jain, and Abbeel 2020) is used to collect a training dataset. Noise $\epsilon$ is gradually added to the original latent $z_{0}$ over $T$ time-steps222 Often, $T$ is chosen to be a large number (e.g., $1000$) such that $z_{T}$ is close to the unit Gaussian noise. , and the info corresponding to a subset of time-steps $t$ is recorded in a form of $\{t,z_{t},\epsilon_{t}\}$. Rombach et al. (2022) trained a UNet (Ronneberger, Fischer, and Brox 2015) denoiser, parameterized by $\theta$, to predict the added noise $\epsilon_{t}$ conditioned on the latent code $z_{t}$, the time-step $t$, and the text embedding $c$ from the text encoder:
$\mathcal{L}_{\text{train}}=\mathbb{E}_{z~{}\sim\mathcal{E}(x),y,\epsilon\sim\mathcal{N}(0,1)}\Big{[}||\epsilon-\epsilon_{\theta}(z_{t},t,c)||^{2}_{2}\Big{]}.$ (1)
@@
To generate novel images conditioned on text prompt, the already-trained UNet is called to sequentially denoise a latent code sampled from the unit Gaussian distribution $z_{T}\sim\mathcal{N}(0,1)$ and conditioned on the text embedding $c$, by traversing back through $T$ time-steps following the RGP. The denoising process results in producing a clean final latent code $z_{0}$, which is passed to the VAE’s decoder to obtain the respective generated image $\mathcal{I}=\mathcal{D}(z_{0})$.
Cross-Attention (CA).
@@
The CA unit is responsible for injecting the prompt information into the model. A CA map at time-step $t$ is denoted by $A_{t}$, having a $P\times P\times N$ tensor size, where $P$ is the spatial dimension of the intermediate feature map and $N$ is the number of prompt tokens. The CA layer weights project the text embedding $c$ into keys $K$ and values $V$, while queries $Q$ are mapped from the intermediate feature map of the UNet. Subsequently, the CA map $A_{t}$ is calculated as $A_{t}=\textit{Softmax}(\frac{QK^{T}}{\sqrt{d}})$.333 As such, $A_{t}[i,j,:]$ defines a probability distribution over the tokens. We represent the CA map for token $n$ by $A_{t}^{n}=A_{t}[:,:,n]$. We utilize the same Gaussian filter $G$ from  (Chefer et al. 2023) to smooth the CA map (i.e., $G(A_{t}^{n})$) for each token. Smoothing the CA maps ensures that the influence of neighboring patches is also reflected in the probability of each patch.
Text Encoder.
@@
Conditioned on the text prompt $y$, the CLIP (Radford et al. 2021) text encoder produces an $N\times H$ text embedding tensor $c^{y}$, where $c^{y}=\text{CLIP}(y)=[c^{y,1},c^{y,2},...,c^{y,N}]$, $N$ denotes the number of tokens, and $H$ is dimension of the text embedding. To satisfy the requirements for the classifier-free guidance (CFG) (Ho and Salimans 2022) approach, an unconditional text embedding is also generated from the null text $\varnothing$, where $c^{\varnothing}=\text{CLIP}(\varnothing)=[c^{\varnothing,1},c^{\varnothing,2},...,c^{\varnothing,N}]$. The conditional token embedding for token $n$ is denoted as $c^{y,n}\in\mathbb{R}^{H}$ and the unconditional token embedding for token $n$ is denoted as $c^{\varnothing,n}\in\mathbb{R}^{H}$.
Prompt Weighting
@@
Prompt weighting (Face 2023a; Stewart 2023) is a popular technique among SD art creators for obtaining more control over the image generation process. Users can manually specify per-token weight coefficients $\phi=[\phi^{1},\phi^{2},...,\phi^{N}]$ to emphasize or de-emphasize the semantics behind each token. Prompt weighting obtains a weighted text embedding $\tilde{c^{y}}$ by interpolating between the conditional text embedding $c^{y}$ and the unconditional text embedding $c^{\varnothing}$ with the weight coefficients $\phi$ as follows:
@@
The SD model adopts the CFG (Ho and Salimans 2022) approach to control the strength of conditional information on the generated image. This strength is controlled with a guidance scale $\beta$. With the prompt weighted text embedding $\tilde{c^{y}}$, CFG updates the UNet’s predicted noise as follows:
@@
The manual prompt weighting technique (Face 2023a; Stewart 2023) uses fixed weight coefficients throughout the entire RGP. Moreover, for every single prompt, the users must manually determine the weight coefficients through expert knowledge.444 The manual prompt weighting library (Stewart 2023) suggests a bounded range of $\phi\in[0.5,1.5]$, since too large or too small of a weighting coefficient has been empirically shown to create low-quality results.
Related works
@@
A recent line of work attempts to automate this task by developing methods for training-free improvement of the semantics compliance of SD, which is an open-sourced SOTA model. In the remainder of this section, we describe the most prominent related works in the literature and touch on their potential limitations.
StructureDiffusion (Feng et al. 2022) used a language model to obtain hierarchical structures from the prompt in the form of noun phrases. They then used the average of the respective CA maps (Hertz et al. 2022) as the output of the CA unit. This method, however, is limited in terms of its ability to generate images that diverge significantly from those produced by SD (Rombach et al. 2022).
Attend & Excite (A&E) (Chefer et al. 2023) introduced the novel concept of GSN, which refers to manipulating the latent codes to guide the RGP towards improving the image alignment with the prompt. They proposed a loss function based on the CA maps to update the latent codes at each time-step. However, their loss function does not explicitly consider the modifier binding issue. Moreover, it only considers a point estimate (i.e., the highest pixel value among all CA maps of noun objects), making it susceptible to having robustness issues.
Divide & Bind (D&B) (Li et al. 2023) adopted the GSN idea (Chefer et al. 2023), but proposed a novel composite loss function that addresses both catastrophic neglect with a Total Variation (TV) loss, as well as the modifier binding issue with a Jensen Shannon Divergence (JSD) loss.
SynGen (Rassin et al. 2023) employed a language model (spaCy’s transformer-based dependency parser (Honnibal and Montani 2017)) to syntactically analyze the prompt. This was to first identify entities and their modifiers, and then use CA maps to create an objective that encourages the modifier binding only. As a result, if the latent code misses an object for some reason, their binding-only loss cannot compensate for this error.
Method

### v2

@@ Abstract
Text-to-image (T2I) diffusion models have demonstrated impressive capabilities in generating high-quality images given a text prompt. However, ensuring the prompt-image alignment remains a considerable challenge, i.e., generating images that faithfully align with the prompt’s semantics. Recent works attempt to improve the faithfulness by optimizing the noisy image latent code, which potentially could cause the latent code to go out-of-distribution and thus produce unrealistic images. In this paper, we propose FRAP, a simple, yet effective approach based on adaptively adjusting the per-token prompt weights to improve prompt-image alignment and authenticity of the generated images. We design an online algorithm to adaptively update each token’s weight coefficient, which is achieved by minimizing a unified objective function that encourages object presence and the binding of object-modifier pairs. Through extensive evaluations, we show FRAP generates images with higher or comparable prompt-image alignment to prompts from complex datasets, while having a lower average latency compared to recent latent code optimization methods, e.g., $4$ seconds faster than recent method D&B on the COCO-Subject dataset. Furthermore, through visual comparisons and evaluation of the CLIP-IQA-Real metric, we show that FRAP not only improves prompt-image alignment but also generates more authentic images with realistic appearances. We also explore combining FRAP with prompt rewriting LLM to recover their degraded prompt-image alignment, where we observe improvements in both prompt-image alignment and image quality. We release the code at the following link: https://github.com/LiyaoJiang1998/FRAP/.
1 Introduction
@@
Recent text-to-image (T2I) diffusion models (Rombach et al., 2022; Chang et al., 2023; Kang et al., 2023) have demonstrated impressive capabilities in synthesizing photo-realistic images given a text prompt. Faithfulness and authenticity are among the most important aspects of assessing the quality of AI-generated content, where faithfulness is evaluated by prompt-image alignment metrics (Radford et al., 2021; Li et al., 2022; Chefer et al., 2023) and authenticity reflects how realistic the image appears and is usually evaluated by image quality assessment metrics (Wang et al., 2023; Wu et al., 2023). While recent T2I diffusion models are capable of generating realistic images with high aesthetic quality, it has been shown (Chefer et al., 2023; Feng et al., 2022; Wang et al., 2022) that the AI-generated images may not always faithfully align with the semantics of the given text prompt. One typical issue that compromises the model’s ability to align with the given prompt is catastrophic neglect (Chefer et al., 2023), where one or more objects are missing in the generated image. Another common alignment issue is incorrect modifier binding (Li et al., 2023), where the modifiers, e.g., color, texture, etc., are either bound to the wrong object mentioned in the prompt or ignored. The generation can also result in an unrealistic mix or hybrid of different objects.
@@
To improve prompt-image alignment, Attend & Excite (A&E) (Chefer et al., 2023) introduces the Generative Semantic Nursing (GSN) method, which iteratively adjusts the latent code throughout the reverse generative process (RGP) by optimizing a loss function defined on cross-attention (CA) maps to improve attention to under-attended objects. Similarly, Divide & Bind (D&B) (Li et al., 2023) uses GSN to improve the binding between objects and modifiers, by designing an optimization objective that aligns the attention of objects and their related modifiers. Although GSN can alleviate the semantic alignment issues, yet the iterative updates to latent code could potentially drive the latent code out-of-distribution (OOD), resulting in less realistic images being generated. In addition, several recent works investigate prompt optimization (Hao et al., 2023; Valerio et al., 2023; Mañas et al., 2024), which aims to improve image aesthetics while trying to maintain the prompt-image alignment, by rewriting the user-provided text prompt with large language model (LLM). Despite improvements in image aesthetic quality, Mañas et al. (2024) show that rewriting prompts primarily for aesthetics enhancement could harm the semantic alignment between the image and prompt. Therefore, it remains a significant challenge to generate images to ensure prompt-image alignment while achieving a realistic appearance and high image authenticity.
@@
In this paper, we introduce Faithful and Realistic Text-to-Image Generation with Adaptive Prompt Weighting (FRAP), which adaptively adjusts the per-token prompt weighting throughout the RGP of a diffusion model to improve the prompt-image alignment while still generating high-quality images with realistic appearance. We utilize the spaCy (Honnibal & Montani, 2017) language parser to identify object tokens and object-modifier relationships from the prompt. Unlike the conventional manual prompt weighting techniques (Hugging Face, 2023a; Stewart, 2023), we design an online optimization algorithm which is performed during inference and updates each token’s weight coefficient in the prompt to adaptively emphasize or de-emphasize certain tokens at different time-steps, by minimizing a unified objective function defined based on the cross-attention maps of the identified object and modifier tokens to strengthen the presence of the prompted object and encourage modifier binding to objects. Since we optimize each prompt token’s weight coefficient, our approach avoids the above-mentioned out-of-distribution issue associated with directly modifying the latent code. Furthermore, we also introduce a novel usage of FRAP by combining it with LLM-based prompt optimizer to achieve outstanding enhancement to both prompt alignment and generation quality.
@@
We extensively evaluate the faithfulness, overall image quality, and image authenticity of FRAP-generated images via prompt-image alignment metrics, image quality assessment metrics, and an image authenticity metric. We make the following observations:
@@
  • We observe that FRAP achieves higher or comparable faithfulness than recent methods on all prompt-image alignment metrics on the Color-Obj-Scene, COCO-Subject, and COCO-Attribute datasets (Li et al., 2023) which contain complex and challenging prompts, while remaining on par with these methods on simple datasets with prompts created from templates. In the meantime, the images generated by FRAP demonstrate significantly higher authenticity with a more realistic appearance in terms of the CLIP-IQA-Real measure (Wang et al., 2023).
@@
  • While improving prompt-image alignment and generating more realistic and authentic images, our method requires lower average inference latency (e.g., $4$ seconds faster than D&B on the COCO-Subject dataset) and lower number of UNet calls by not making repeated calls of UNet at each time-step during inference, in contrast to the recent GSN methods (Chefer et al., 2023; Li et al., 2023) which calls the UNet multiple times in a refinement loop at each time-step, especially on datasets with complex prompts that require many repeated calls.
@@
  • We observe that although the LLM prompt rewriting method Promptist (Hao et al., 2023) improves the image aesthetic quality, it could harm the semantic alignment between the image and prompt. Moreover, our adaptive prompt weighting approach can be easily integrated with recent prompt rewrite methods (Hao et al., 2023; Valerio et al., 2023; Mañas et al., 2024). We explore applying FRAP to the rewritten prompts of Promptist to recover their degraded prompt-image alignment, where we observe improvements in both the prompt-image alignment and overall image quality.
2 Preliminaries
2.1 Stable Diffusion (SD)
@@
We implement our approach based on the open-source T2I latent diffusion model SD (Rombach et al., 2022). Unlike traditional diffusion models (such as Sohl-Dickstein et al. (2015); Ho et al. (2020); Song et al. (2020)) that work in the pixel space, SD operates in an autoencoder’s latent space, which significantly reduces the amount of compute resources required. SD is composed of several components, including a Variational Auto-Encoder (VAE) (Kingma & Welling, 2013; Rezende et al., 2014)111 The VAE encoder $\mathcal{E}$ maps an input image $x$ into its spatial latent representation $z=\mathcal{E}(x)$, while the decoder $\mathcal{D}$ attempts to reconstruct the same image from $z$ such that $\mathcal{D}\big{(}\mathcal{E}(x)\big{)}\approx x$. , a UNet (Ronneberger et al., 2015) that operates in the latent space and serves as the backbone for image generation, and a Text Encoder that is used to obtain embeddings from the text prompt.
2.1.1 UNet
@@
For training the UNet, the forward diffusion process (Ho et al., 2020) is used to collect a training dataset. Noise $\epsilon$ is gradually added to the original latent $z_{0}$ over $T+1$ time-steps222 Often, $T$ is chosen to be a large number (e.g., $1000$) such that $z_{T}$ is close to the unit Gaussian noise. , and the info corresponding to a subset of time-steps $t$ is recorded in a form of $\{t,z_{t},\epsilon_{t}\}$. Rombach et al. (2022) trained a UNet (Ronneberger et al., 2015) denoiser, parameterized by $\theta$, to predict the added noise $\epsilon_{t}$ conditioned on the latent code $z_{t}$, the time-step $t$, and the text embedding $c$ from the text encoder:
$\mathcal{L}_{\text{train}}=\mathbb{E}_{z~{}\sim~{}\mathcal{E}(x),~{}y,~{}\epsilon~{}\sim~{}\mathcal{N}(\mathbf{0},\mathbf{I})}\Big{[}||\epsilon-\epsilon_{\theta}(z_{t},t,c)||^{2}_{2}\Big{]}.$ (1)
@@
To generate novel images conditioned on text prompt, the already-trained UNet is called to sequentially denoise a latent code sampled from the unit Gaussian distribution $z_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and conditioned on the text embedding $c$, by traversing back through $T+1$ time-steps following the RGP. The denoising process results in producing a clean final latent code $z_{0}$, which is passed to the VAE’s decoder to obtain the respective generated image $\mathcal{I}=\mathcal{D}(z_{0})$.
2.1.2 Cross-Attention (CA)
@@
The CA unit is responsible for injecting the prompt information into the model. A CA map at time-step $t$ is denoted by $A_{t}$, having a $L\times L\times N$ tensor size, where $L$ is the spatial dimension of the intermediate feature map and $N$ is the number of prompt tokens. The CA layer weights project the text embedding $c$ into keys $K$ and values $V$, while queries $Q$ are mapped from the intermediate feature map of the UNet. Subsequently, the CA map $A_{t}$ is calculated as $A_{t}=\textit{Softmax}(\frac{QK^{T}}{\sqrt{d}})$.333 As such, $A_{t}[i,j,:]$ defines a probability distribution over the tokens. We represent the CA map for token $n$ by $A_{t}^{n}=A_{t}[:,:,n]$. We utilize the same Gaussian filter $G$ from Chefer et al. (2023) to smooth the CA map (i.e., $G(A_{t}^{n})$) for each token. Smoothing the CA maps ensures that the influence of neighboring patches is also reflected in the probability of each patch.
2.1.3 Text Encoder
@@
Conditioned on the text prompt $y$, the CLIP (Radford et al., 2021) text encoder produces an $N\times H$ text embedding tensor $c^{y}$, where $c^{y}=\text{CLIP}(y)=[c^{y,1},c^{y,2},...,c^{y,N}]$, $N$ denotes the number of tokens, and $H$ is the dimension of the text embedding. To satisfy the requirements for the classifier-free guidance (CFG) (Ho & Salimans, 2022) approach, an unconditional text embedding is also generated from the null text $\varnothing$, where $c^{\varnothing}=\text{CLIP}(\varnothing)=[c^{\varnothing,1},c^{\varnothing,2},...,c^{\varnothing,N}]$. The conditional token embedding for token $n$ is denoted as $c^{y,n}\in\mathbb{R}^{H}$ and the unconditional token embedding for token $n$ is denoted as $c^{\varnothing,n}\in\mathbb{R}^{H}$.
2.2 Prompt Weighting
@@
Prompt weighting (Hugging Face, 2023a; Stewart, 2023) is a popular technique among SD art creators for obtaining more control over the image generation process. Users can manually specify per-token weight coefficients $\phi=[\phi^{1},\phi^{2},...,\phi^{N}]$ to emphasize or de-emphasize the semantics behind each token. Prompt weighting obtains a weighted text embedding $\tilde{c^{y}}$ by interpolating between the conditional text embedding $c^{y}$ and the unconditional text embedding $c^{\varnothing}$ with the weight coefficients $\phi$ as follows:
@@
The SD model adopts the CFG (Ho & Salimans, 2022) approach to control the strength of conditional information on the generated image. This strength is controlled with a guidance scale $\beta$. With the prompt weighted text embedding $\tilde{c^{y}}$, CFG updates the UNet’s predicted noise as follows:
@@
The manual prompt weighting technique (Hugging Face, 2023a; Stewart, 2023) uses fixed weight coefficients throughout the entire RGP. Moreover, for every single prompt, the users must manually determine the weight coefficients through expert knowledge.444 The manual prompt weighting library (Stewart, 2023) suggests a bounded range of $\phi\in[0.5,1.5]$, since too large or too small of a weighting coefficient has been empirically shown to create low-quality results.
3 Related works
@@
A recent line of work attempts to automate this task by developing methods for training-free improvement of the semantics compliance of SD, which is an open-sourced text-to-image generation diffusion model. In the remainder of this section, we describe the most prominent related works in the literature and touch on their potential limitations.
StructureDiffusion (Feng et al., 2022) used a language model to obtain hierarchical structures from the prompt in the form of noun phrases. They then used the average of the respective CA maps (Hertz et al., 2022) as the output of the CA unit. This method, however, is limited in terms of its ability to generate images that diverge significantly from those produced by SD (Rombach et al., 2022).
Attend & Excite (A&E) (Chefer et al., 2023) introduced the novel concept of GSN, which refers to manipulating the latent codes to guide the RGP towards improving the image alignment with the prompt. They proposed a loss function based on the CA maps to update the latent codes at each time-step. However, their loss function does not explicitly consider the modifier binding issue. Moreover, it only considers a point estimate (i.e., the highest pixel value among all CA maps of noun objects), making it susceptible to having robustness issues.
Divide & Bind (D&B) (Li et al., 2023) adopted the GSN idea (Chefer et al., 2023), but proposed a novel composite loss function that addresses both catastrophic neglect with a Total Variation (TV) loss, as well as the modifier binding issue with a Jensen Shannon Divergence (JSD) loss.
SynGen (Rassin et al., 2023) employed a language model (spaCy’s transformer-based dependency parser (Honnibal & Montani, 2017)) to syntactically analyze the prompt. This was to first identify entities and their modifiers, and then use CA maps to create an objective that encourages the modifier binding only. As a result, if the latent code misses the generation of an object, their binding-only loss cannot compensate for this error.
4 Method